a solid of constant density is bounded below by the plane z = 0 , on the sides by the ellptical cylinder x^2 + 4y^2 = 4, and above by the plane z = 2 - x . calculate a) volume b) center of mass c) moment of inertial about x-axis , y-axis and z-axis


1
Expert's answer
2021-03-31T11:06:25-0400

a) The volume

"V=\iiint dxdydz"

Now, "z=0;x^2+4y^2=4;z=2-x"

Then, "\text{\textbraceleft}0\eqslantless z \eqslantless2-x, 0\eqslantless x \eqslantless \sqrt{4-4y^2}, 0\eqslantless y\eqslantless1 \text{\textbraceright}"

So the volume integral becomes

"V=\int_0^1 \int_0^{\sqrt{4-4y^2}}\int_0^{2-x}dzdxdy"

"V=\int_0^1 \int_0^{\sqrt{4-4y^2}}(2-x)dxdy"

"V=\int_0^1 [2x- \frac{x^2}{2}]_0^{\sqrt{4-4y^2}}dy"

"V=2[ \frac{y^3}{3}+y \sqrt{1-y^2}-y+arcsin (y)]_0^1"

"V=\pi- \frac{4}{3}"


b) Center of mass 

"M=\int_{-2}^2 \int_{-0.5\sqrt{4-4y^2}}^{0.5\sqrt{4-4y^2}}\int_0^{2-x}dzdydx"

"M=\int_{-2}^2 \int_{-0.5\sqrt{4-4y^2}}^{0.5\sqrt{4-4y^2}}[2-x]dydx"

"M=\int_{-2}^2 ((2-x)\sqrt{4-4y^2})dx"

"M=\int_{0}^2 2\sqrt{4-4y^2}-x\sqrt{4-4y^2})dx=4 \pi"


"M_{yz}=\int_{-2}^2 \int_{-0.5\sqrt{4-4y^2}}^{0.5\sqrt{4-4y^2}}\int_0^{2-x}xdzdydx"

"M_{yz}=\int_{-2}^2 \int_{-0.5\sqrt{4-4y^2}}^{0.5\sqrt{4-4y^2}}x[2-x]dydx"

"M_{yz}=\int_{-2}^2 ((2x-x^2)\sqrt{4-4y^2})dx"

"M_{yz}=-2\int_{0}^2 x^2\sqrt{4-4x^2} dx"

Let "x= 2sin \theta; dx=2cos \theta d\theta"

"M_{yz}=-32 \int_0 ^{\frac{\pi}{2}}sin^2 \theta cos^2 \theta d \theta =-2 \pi"


"M_{xz}=\int_{-2}^2 \int_{-0.5\sqrt{4-4y^2}}^{0.5\sqrt{4-4y^2}}\int_0^{2-x}ydzdydx"

"M_{xz}=\int_{-2}^2 \int_{-0.5\sqrt{4-4y^2}}^{0.5\sqrt{4-4y^2}}y[2-x]dydx"

"M_{xz}=\frac{2-x}{2}[\frac{(\sqrt{4-4y^2})^2}{2}-\frac{(\sqrt{4-4y^2})^2}{2}] =0"


So, the center of mass is

"\bar{x}= \frac{M_{yz}}{M}= \frac{-2 \pi}{4 \pi}= -\frac{1}{2}"

"\bar{y}= \frac{M_{xz}}{M}= \frac{0}{4 \pi}= 0"


c) Moment of inertial about the x-axis, y-axis, and z-axis

"M_{yz}=\int_{-2}^2 \int_{-0.5\sqrt{4-4y^2}}^{0.5\sqrt{4-4y^2}}\int_0^{2-x}xdzdydx"

"M_{yz}=\int_{-2}^2 \int_{-0.5\sqrt{4-4y^2}}^{0.5\sqrt{4-4y^2}}x[2-x]dydx"

"M_{yz}=\int_{-2}^2 ((2x-x^2)\sqrt{4-4y^2})dx"

"M_{yz}=-2\int_{0}^2 x^2\sqrt{4-4x^2} dx"

Let "x= 2sin \theta; dx=2cos \theta d\theta"

"M_{yz}=-32 \int_0 ^{\frac{\pi}{2}}sin^2 \theta cos^2 \theta d \theta =-2 \pi"


"M_{xz}=\int_{-2}^2 \int_{-0.5\sqrt{4-4y^2}}^{0.5\sqrt{4-4y^2}}\int_0^{2-x}ydzdydx"

"M_{xz}=\int_{-2}^2 \int_{-0.5\sqrt{4-4y^2}}^{0.5\sqrt{4-4y^2}}y[2-x]dydx"

"M_{xz}=\frac{2-x}{2}[\frac{(\sqrt{4-4y^2})^2}{2}-\frac{(\sqrt{4-4y^2})^2}{2}] =0"


"M_{xy}=\int_{-2}^2 \int_{-0.5\sqrt{4-4y^2}}^{0.5\sqrt{4-4y^2}}\int_0^{2-x}zdzdydx"

"M_{xy}=\int_{-2}^2 \int_{-0.5\sqrt{4-4y^2}}^{0.5\sqrt{4-4y^2}}z[2-x]dydx"

"M_{xy}=\int_{-2}^2 \sqrt{4-x^2})dx+\frac{1}{4}[x^2\int_{-2}^2 \sqrt{4-x^2})dx-\int_{-2}^2 \sqrt{4-x^2})dx]"

"M_{xy}=-2\int_{0}^2 x^2\sqrt{4-4x^2} dx"

"M_{xy}=2* \frac{\pi}{2}+4 \pi =5 \pi"